LWS Upgrade Project News
updated 08 August 98
The LWS upgrade project was scheduled for a "readiness review" on 18Aug98.
The most recent goal was to install LWS on the Keck I telescope and begin
commissioning on the night sky in early September. However, the science
grade device turned out to have a major problem, an instability of the
bias voltage caused by a breakdown of the guard ring. The review
and commissioning are being postponed until the detector can be replaced.
The new date for the review is not known at this time. Boeing, the detector
manufacturer has been very responsive and cooperative in diagnosing
the problem. They have agreed to replace the device with one thoroughly
tested and proven not to have this problem. It is hoped that the
schedule setback will not be too severe and LWS commissioning will proceed
in the fall of 1998.
scheduled for 19Aug98
Detector installation complete, cool down cycle and testing to begin
scheduled for 16Aug98
Replacement science grade detector scheduled to arrive in Hawaii
Telecon occurs with members of LWS team and Boeing personnel.
Determined that most likely cause of problem is breakdown of detector guard
ring leading to a drop in bias voltage.
Boeing to replace with a device that is shown to not have the guard ring
Detector tests continue with detector in good state.
System noise shown to be much better but not quite background limited
System noise shown to be a function of chop frequency, or 1/f noise.
Detector proves to be very good when not experiencing repsonsivity problem.
First Cold Tests of science grade device.
Problem discovered: instability in signal level, evident on image and on
Most evident with flux on chip and temperature controlled to greater 8degK
Problem corrects itself when power is cycled to
26Jul98 Science grade device ,Boeing 128^2 As:Si BIB Moderate
Flux, arrives in Hawaii
System testing with Engineering array
System Noise problem improved significantly to around 0.7 adu
Stray radiation problem significantly improved, now negligible for typical
First cold tests () of the engineering grade chip were successful!
One dead channel on engineering array.
Sensitive for all filters, sensitivity not yet quantified.
Problem discovered: system/detector noise being higher than background
Stray radiation in dewar detected, limiting ability to test and quantify
Temperature control of detector worked very well in range 6 to 11 deg K
Project review and summit commissioning anticipated for late summer 98
LWS Image of Test Aperture at 11.7 microns, 30May98